OpenStudy (anonymous):

fifth term of a geometric sequence is 112 and the ninth term is 1,792. Find the common ratio of the sequence.

OpenStudy (alekos):

For a geometric seq An+1/An = r the common ratio and An+2/An = r^2 and so on... what do you think An+4/An will equal?

OpenStudy (anonymous):

 use this formula where: an = the nth therm a1 = the first term n = the number of term r = the ratio between terms